A certian plane can travel 1500 miles in 5 hours when flying into the wind. The same plane can travel 2000 miles in 5 hours when flying with the wind. Find the rate of the wind.
Found 2 solutions by mananth, Alan3354:
Answer by mananth(16949)

A certian plane can travel 1500 miles in 5 hours when flying into the wind. The same plane can travel 2000 miles in 5 hours when flying with the wind. Find the rate of the wind.
let speed of wind be xmph
speed of plane in still air be y
distance = rate * time
1500 = 5(y-x)
similarly 2000 = 5(y+x)
1500 = 5y-5x
2000 = 5y+5x
Add the two equations
weget 10 y= 3500
y= 350 mph
.
1500 = 5y-5x
plug the value of y in the equation
1500 = 1750-5x
-5x= -250
x= 50 mph the speed of the wind


Answer by Alan3354(69443)

A certian plane can travel 1500 miles in 5 hours when flying into the wind. The same plane can travel 2000 miles in 5 hours when flying with the wind. Find the rate of the wind.
-----------------------
1500/5 = 300 mph
2000/5 = 400 mph
The plane's airspeed is the average = (300+400)/2 = 350 mph
The wind speed is the difference, = 400-350 = 50 mph